OpenStudy (anonymous):

.....!Help Please!..... Generate the first 5 terms of this sequence: f(1) = 2 and f(2) = 3, f(n) = f(n - 1) + f(n - 2), for n > 2. 2, 3, 5, 7, 9 2, 3, 4, 7, 11 2, 3, 5, 9, 11 2, 3, 5, 8, 13

OpenStudy (anonymous):

yo person

OpenStudy (anonymous):

Hey! ^.^

OpenStudy (anonymous):

i dont know if i can solve this but let me look is that all right?

OpenStudy (anonymous):

Okay! that's fine! Thanks for helping! :)

OpenStudy (anonymous):

the answer is the last one 2,3,5,8 and 13

OpenStudy (anonymous):

(i think)

OpenStudy (anonymous):

Ok :) Can you explain to me why you think that's the answer so I can understand? ^.^

OpenStudy (anonymous):

sure id ont know if this will help though. like this put values of n and calculate n=3 f(3)=f(3-1)+f(3-2)=f(2)+f(1)=3=2=5

OpenStudy (anonymous):

Okay! Thank you so much! @zendaya-replay-guy ^.^

OpenStudy (anonymous):

u r welcome